OpenStudy (anonymous):

A tank containing oil is in the shape of a downward-pointing cone with its vertical axis perpendicular to ground level. (See a graph of the tank here.) In this exercise we will assume that the height of the tank is 10 ft , the circular top of the tank has radius 5 ft, and that the oil inside the tank weighs 56 lb per cubic ft. How much work does it take to pump oil from the tank to an outlet that is 3 ft above the top of the tank if, prior to pumping, there is only a half-tank of oil? Note: "half-tank" means half the volume in the tank.

OpenStudy (anonymous):

Do you know any physical laws, equations that could help? @skimo16

OpenStudy (anonymous):

I thought this might be a pressure problem, but it looks like it's more of just a work problem. So I suppose the only equation of pertinence is: \[ \vec{W} = \int \vec{F} \cdot d\vec{r} \]

OpenStudy (anonymous):

So we need to come up with some equations of force here, with respect to distance.

OpenStudy (anonymous):

"56 lb per cubic ft" is our Force/Volume, so we need an equation for volume now.

OpenStudy (anonymous):



OpenStudy (anonymous):

We'll integrate with respect to \(x\).

OpenStudy (anonymous):

The volume at any \(x\) value will be: \(\pi r(x)^2dx\) where \(r(x)\) is our radius in terms of \(x\) and \(dx\) is our thickness

OpenStudy (anonymous):

The distance it must travel for any \(x\) is just \(x\), so our work for each is thus: \(x \times (56\pi r(x)^2dx)\)

OpenStudy (anonymous):

All together we have: \[ \Large \int_{x_i}^{x_f} 56xr(x)^2dx \] So what we need is \(r(x) \) and our limits of integration \(x_i \ x_f\).

OpenStudy (anonymous):

\[ \Large \int_{x_i}^{x_f} 56x\pi r(x)^2dx \]

OpenStudy (anonymous):

@skimo16 Is it starting to make sense?

OpenStudy (anonymous):